Identify the critical value of each given problem. Find the rejection region and sketch the curve on a separate sheet of paper.

“According to the radio announcer, the average price of kilogram of pork liempo is more than ₱210.00. However, a sample of 15 prices randomly collected from different markets showed an average of ₱215.00 and standard deviation of ₱9.00. Using 0.05 level of significance, is there sufficient evidence to conclude that the average price of pork liempo is more than ₱210.00?”



1
Expert's answer
2022-05-31T17:43:32-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le210"

"H_1:\mu>210"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=14" and the critical value for a right-tailed test is "t_c = 1.76131."

The rejection region for this right-tailed test is "R = \{t:t> 1.76131\}."


The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{215-210}{9/\sqrt{15}}=2.1517"


Since it is observed that "t=2.1517> 1.76131=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=14" degrees of freedom, "t=2.1517" is "p= 0.02468," and since "p= 0.02468<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 210, at the "\alpha = 0.05" significance level.